OpenStudy (anonymous):

The Coefficient of kinetic friction between the tires of your car and the roadway is "μ". (a) If your initial speed is "v" and you lock your tires during breaking, how far do you skid? (b) What is the stopping distance for a truck with twice the mass of you car, assuming same initial speed and coefficient of kinetic friction. I'm stuck on (b), my mass cancels out every time, but that doesn't seem right

OpenStudy (anonymous):

Can u show wat u have done

OpenStudy (noelgreco):

If you answered (a) correctly the answer would be clear.

OpenStudy (anonymous):

F(friction) / 2m = a (new) v(final)^2 = v(inital)^2 + 2ad substitute the new a into the second equation assume final velocity is 0 since it's stopped